Consider the demand function for processed pork in Canada -254.00-10p+20p+3pc+0.002Y The supply function for processed pork in Canada is: -274.00+36p-60Pm p is the price of pork Q is the quantity of pork demanded Pb is the price of beef $4 per kg Pc is the price of chicken $3 per kg (measured in millions of kg per year) Y is the income of consumers $12,500 of kg per year) Yis Ph is the price of a hog-$1.50 per kg Solve for the equilibrium price and quantity for pork The equilibrium price of pork i nd the equilibrium quantity of pork ismillion kg per year. (Enter numeric responses using real numbers rounded up to two decimal places.)
Answer #1

Demand equation

Qd = 254-10p+20pb+3pc+0.002Y

Substitute the given values in the demand equation

Qd = 254-10p+20(4)+3(3)+0.002(12500)

Qd= 254-10p+80+9+25

Qd=368-10p

Supply equation

Qs = 274+36p-60ph

Substitute the given values in the supply equation

Qs = 274+36p-60(1.50)

Qs = 274+36p-90

Qs=184+36p

At market equilibrium, Demand = Supply

Qd = Qs

368-10p = 184+36p

368-184 = 36p+10p

184=46p

P= 184/46

P = 4

Equilibrium price of pork is $4

Substitute P = $4 in Qs

Qs=184+36p

Qs = 184 + 36(4)

Q = 184+144

Q = 328

Equilibrium quantity of pork = 328 million kg per year